OpenStudy (anonymous):

Solve the equation, if possible. Give answer in radical form. x^2=48

hartnn (hartnn):

x^2 = 48 take square root on both sides, what u get ?

OpenStudy (anonymous):

6.93

hartnn (hartnn):

lets not use any calculator. x^2 = 48 \(x = \pm \sqrt {48}\) can you factor 48 ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so, 48 =... ? (please factor it)

OpenStudy (anonymous):

so i got 2 pairs of 2s and only one 3

hartnn (hartnn):

correct! \(\sqrt {48} = \sqrt {(2 \times 2) (2 \times 2) 3}\)

hartnn (hartnn):

so, those 2 pairs can be taken out of the square root sign \(\sqrt {48} = 2 \times 2 \times \sqrt3\)

hartnn (hartnn):

finally, \(\sqrt {48} = 4 \sqrt 3\) thats it!

OpenStudy (anonymous):

So you just times the pair???

hartnn (hartnn):

\(x = + 4 \sqrt 3 \\ x = -4 \sqrt 3\) yes, we use , \(\sqrt {a \times a} = a \)

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^